OpenStudy (anonymous):

Help Please Graph the exponential function. y = 3(1.9)x

OpenStudy (anonymous):

http://prntscr.com/12cfh3

OpenStudy (anonymous):

@.Sam.

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

is it y= 3* 1.9^x ??

hartnn (hartnn):

look for x=0 that is the point where graph cuts y=axis from equation, y= 3*1.9^0 = ... ?

OpenStudy (anonymous):

wait

hartnn (hartnn):

which one is C ? the middle one ?

OpenStudy (anonymous):

sorry i asked the wrong question this was the question i intended to ask Graph the exponential function. 4(3)^x

OpenStudy (anonymous):

http://prntscr.com/12cglw

OpenStudy (anonymous):

http://prntscr.com/12cgnf

OpenStudy (anonymous):

I sit B

hartnn (hartnn):

when x increases, the value of 4 (3)^x also increases. also, at x=0, y= 4(3)^0 = 4 from these you can say that its B, yes.

OpenStudy (anonymous):

http://prntscr.com/12cgwg

OpenStudy (anonymous):

http://prntscr.com/12cgxw

OpenStudy (anonymous):

Is it A

hartnn (hartnn):

st, x=0, y=... ?

OpenStudy (anonymous):

This will be graphed http://prntscr.com/12ch4y

hartnn (hartnn):

yes, that graph iss correct.

OpenStudy (anonymous):

option A looks similar

hartnn (hartnn):

yeah, it looks similar, but the intersection there is around y= 3 instead of y= 0.2 or 1/5

hartnn (hartnn):

its not B or D for sure.

OpenStudy (anonymous):

http://prntscr.com/12ci0r

OpenStudy (anonymous):

about this graph

hartnn (hartnn):

yes, its that graph.

OpenStudy (anonymous):

but it was marked

OpenStudy (anonymous):

wrong

hartnn (hartnn):

hmm...one thing i know for sure, is that graph is correct. there must be mistake in the system .... ?